OpenStudy (anonymous):

How can we create equations in one variable and use them to solve problems? How can we solve linear equations in one variable?

OpenStudy (anonymous):

linear equations in 1 variable has one variable x which can be written as ax+b=0

OpenStudy (anonymous):

example: x+5=-3 x/3+1/2= 2/3

OpenStudy (anonymous):

wow thank you!! you know question number 1?

OpenStudy (anonymous):

your welcome ) which question?

OpenStudy (anonymous):

How can we create equations in one variable and use them to solve problems? i don't know this :(

OpenStudy (anonymous):

i will send you the link juts have a look.. https://learnzillion.com/lessons/763-create-linear-equations-to-solve-problems

OpenStudy (anonymous):

ohh okey thank you!! you are my lifesaver haha

OpenStudy (anonymous):

no problem)